OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2), Q(8, -4)

OpenStudy (marigirl):

First we need to workout our gradient or m value m=y2-y1/x2-x1 m=(-4-2)/(8-6) m=(-6)/(2) m=-3 next we can form the equation by using the formula y-y1=m(x-x1) y-2=-3(x-6) y-2=-3x+18 y=-3x+20 (gradient and intercept form) y+3x=20(standard form)